40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5277     Entry time: Mon Aug 22 16:28:44 2011
Author: Suresh 
Type: Update 
Category: General 
Subject: SRM tower shifted on Friday 

[Kiwamu, Suresh]   This is a belated elog entry from last Friday night + Saturday morning! 

    We shifted the SRM tower across the beam and away from the door by about 5mm. 

                  After the input beam from MC was aligned to the Y-arm,  Kiwamu noticed that the AS beam was being clipped and that the correction had to start from SRM onwards as the beam had become offcentered on the SRM.  So we shifted the SRM tower by about 5mm away from the door and transverse to the beam and rotated it by a few degrees to center the OSEM offsets.  After this we aligned all optics along the AS beam path to extract the AS beam from the chamber.    We then aligned each optic in the vertex so that their beams overlap at the AS port with the reflection from ITMY.  Then we aligned BS to center the beam on ETMX and then looked for flashes from the Y arm.

      At this point Kiwamu checked and found that the input beam from the MC had shifted.  It was landing on the ITMY about 5mm below the center.  And (inexplicably) it was still centered on ETMY!  The Y- green which traced the cavity axis (since this was still flashing) was not coincident with the IR beam.   So all the work we did in aligning the AS beam and the vertex optics work was lost..... and had to be done again.

 

 

ELOG V3.1.3-